OpenStudy (lilsis76):

can someone help me... 2 Al + 6 HCL <=> 3 H + 2 AlCl How many mL of 6.0M hydrochloric acid must react with excess aluminum to form 0.375 moles of aluminum chloride?

OpenStudy (lilsis76):

I took chemistry last year and I cant remember the steps to take to solve this answer.

OpenStudy (aaronq):

Find the amount of moles of HCl you need, first, then convert that to volume using: Molarity = moles of solute/L of solution To find the amount of moles you need to use their stoichiometric ratio: for a general equation, where lower case are the stoichiometric coefficients and upper case are the species: aA + bB <-> cC + dD moles of A/a=moles of B/b from this you can find how many moles of HCl you need.

OpenStudy (lilsis76):

@aaronq. OKay let me try. Im going to look up the masses on the periodic table first. Im just telling you my steps as I go to see if I am doing it correct. I feel so completly lost beause Its been awhile since I did conversions

OpenStudy (aaronq):

you don't need the molar masses.

OpenStudy (lilsis76):

you dont? Hmm...okay then well i have: 6HCl in the equation then. how many moles mL of 6.0M HCl must react with excess aluminum to form .375 moles of Aluminum chloride. okay so the 6.0M HCL will be like this \[\frac{ 6.0 m }{ L }\] right?

OpenStudy (aaronq):

nope. "M" means molarity, so 6.0 M means that there are 6 moles of solute in 1 L of water. i'll help you get started, what you need is "0.375 moles of AlCl3". use the stoichiometric coefficients from the balanced equation, V V "2 Al + 6 HCL <=> 3 H + 2 AlCl3" (moles of HCl)/6= (0.375 moles of AlCl3)/2 solve for moles, then use those moles in this equation: Molarity=moles of solute/L of solution 6.0 M = moles (you found in previous part)/L of solution you're finding L, don't forget to convert to mL

OpenStudy (lilsis76):

haha im back. said connection was lost @aaronq

OpenStudy (aaronq):

thats okay, did you understand what i wrote?

OpenStudy (lilsis76):

im reading it now. im going step by step. Sorry if im slow, i completly lost all chem from my brain haha

OpenStudy (aaronq):

lol i know, use it or lose it. let me know if you need further help

OpenStudy (lilsis76):

okay to convert to mL I multiply by 1000?

OpenStudy (lilsis76):

WAIT! I cross multiply?

OpenStudy (aaronq):

divide by 1000 1 mL = 0.001 L

OpenStudy (lilsis76):

okay then, hold on its.....here ill draw it if i can

OpenStudy (lilsis76):

 is this right so far? @aaronq